OpenStudy (anonymous):

A line with a slope of 3/14 passes through points (7,3k) and (0,k), What is the value of k?

OpenStudy (anonymous):

Find the slope of (7,3k) and (0,k),

OpenStudy (anonymous):

\[m = \frac{ (y2-y1) }{ (x2-x1) }\] \[m = \frac{ 3 }{ 14 }\]

OpenStudy (anonymous):

Establish a system of equations

OpenStudy (anonymous):

Use what yahoo said: m = (y2 - y1) / (x2 - x1)

OpenStudy (anonymous):

( x1, y1) = ( 7 , 3k)

OpenStudy (anonymous):

(x2 . y2) = ( 0 , k)

OpenStudy (anonymous):

Replace at equation

OpenStudy (anonymous):

the question said m=3/14

OpenStudy (anonymous):

3/14 = ( k - 3k) / (0 - 7) 3/14 = -2k / -7

OpenStudy (anonymous):

7 (3) / 14 = 2k 3/4 = k

OpenStudy (anonymous):

well It was a pleasure to help you Saludos desde Perú